# More Tokens, Lower Precision: Towards the Optimal Token-Precision Trade-off in KV Cache Compression

### v1

@@ 6.2 Scaling Effect on Quantized Pruning
Figure 4: Scaling effect on Llama family models, with PyramidKV-based quantized pruning. All models are under fixed ratio of KV cache budget.
To investigate the impact of model scaling on quantized pruning, we conducted experiments on three models from the Llama series: Llama3-8B, Llama3.2-3B, and Llama3.2-1B. For both the Base models and Instruct models, we evaluated their performance on LongBench under two fixed KV cache budgets: 1/16 and 1/64.
As shown in Figure 4, we observe that quantized pruning consistently achieves better performance across all scaling levels. The performance gap between quantized pruning and standalone KV pruning methods remains relatively stable across different model scales. Notably, when the KV cache budget is relative small to 1/64 , the performance improvement brought by quantized pruning is higher compared to 1/16 KV cache budget, which aligns with the conclusions we observed earlier in Q2. For Base models, although the performance improvement from scaling is smaller compared to Instruct models, quantized pruning still provides a noticeable performance boost.
@@ 6.3 Ablation on Quantization Strategies
While there has been extensive research on strategies for KV cache quantization, it remains unclear whether existing quantization strategies remain effective when combined with KV pruning methods. In this section, we aim to investigate the impact of KV quantization strategies and group size on quantized pruning, and present our results in Figure 5 and Table 3.
Quantization methods
We explore the methods in FlexGen (Sheng et al., 2023), KIVI (Liu et al., 2024c), and KVQuant (Hooper et al., 2024). To elaborate, for the FlexGen methods, KV quantization is applied to both the key and value caches along the token dimension, grouping every 64 elements without filtering outlier numbers. We modify the FlexGen by (1) filtering 1% of outlier numbers in both the key and value caches, as mentioned in KVQuant (2) quantizing the key along the channel dimension, as in KIVI and (3) combining (1) and (2). These correspond to the results labeled as FlexGen+Outlier 1%, KIVI, and KIVI+Outlier 1% in the Figure 5.
We can observe that none of the quantization strategies show significant performance degradation when combined with KV pruning methods, demonstrating the relative stability of quantized pruning. The KIVI method consistently outperforms FlexGen across various models and KV pruning methods. The improvement is particularly pronounced for PyramidKV on the Mistral model, underscoring the significance of quantizing key states along the channel dimension. Filtering 1% of outlier numbers proves effective for the FlexGen strategy but yields limited improvements for KIVI. It shows some benefit on Llama models but can result in negative gains on the Mistral model.
Overall, KIVI demonstrates strong performance when combined with KV pruning methods, while other KV quantization strategies also maintain good results, highlighting the stability of quantized pruning.
Group Size
We then analyzed the impact of group size during KV quantization. Using the SnapKV and PyramidKV methods to retain 512 tokens, we experimented with 4-bit quantization and observed the performance variations when the group sizes were set to 32, 64 and 128.
Model Method GroupSize
32 64 128
Llama-3 SnapKV 40.42 39.55 38.87
PyramidKV 40.33 39.65 38.91
Mistral-v0.2 SnapKV 40.45 40.30 40.09
PyramidKV 40.31 40.49 40.03
Table 3: The impact of group size for quantized pruning on LongBench, remaining 512 KV tokens in 4-bit.
As shown in Table 3, smaller group sizes lead to performance improvements at the cost of higher memory usage. Reducing the group size from 128 to 64 resulted in a notable improvement, but further decreasing it from 64 to 32 yielded minimal gains for the Mistral model. Therefore, we set the default quantization group size to 64 to balance performance and memory usage in our experiments.
6.4 Exploration on Layer-Wise Quantized Pruning
Inspired by the observation that different layers may have varying requirements for the number of tokens in PyramidKV (Cai et al., 2024) and PyramidInfer (Yang et al., 2024a), we further investigate whether the demands for precision and preserved tokens are consistent across layers. To explore this, we use the best-performing configuration from previous experiments, 4-bit precision with 4×4$\times$4 × tokens, as the baseline and compare it against layer-wise configurations adopting 8-bit precision with 2×2$\times$2 × tokens and 16-bit precision with 1×1$\times$1 × tokens. Using SnapKV as the KV pruning method, we present the results for Llama-3 and Mistral-v0.2 under two budget constraints in the Figure 6. Configurations are modified every 4 layers for the initial and final layers, while intermediate layers are reconfigured every 8 layers. The x-axis indicates the layers where the modified configurations are applied, while the y-axis shows the relative change to the baseline (4-bit precision with 4×4$\times$4 × tokens) on LongBench and RULER-4k.
@@ 6.4 Exploration on Layer-Wise Quantized Pruning
Figure 6: The results of layer-wise quantized pruning on Llama-3-8B-Instruct, with SnapKV as pruning method. We use 4×4$\times$4 × KV token 4-bit as baseline and report the relative change.
@@
Initially, it is evident that for most layers, transitioning from 4×4$\times$4 × tokens with 4-bit precision to higher precision and fewer tokens results in a performance decline under constrained KV cache budgets. Specifically, the shift to 8-bit shows a relatively minor performance drop, whereas moving to 16-bit with fewer preserved tokens leads to a more significant decrease. These layers-wise trade-off conclusions are consistent with our experiments before.
@@
Notably, modifying intermediate layers causes a drastic performance decline, while adjustments made at the initial and final layers result in comparatively smaller performance reductions. This effect is especially pronounced in retrieval-related tasks such as RULER-4k, where significant performance differences are observed. On LongBench, changes are less significant, with a notable performance drop only observed at 16-bit precision. These findings highlight that, under the same memory budget, preserving more tokens in intermediate layers is crucial for the performance, while the token-precision trade-off in the initial and final layers exerts a more balanced influence on the results.
@@ 7 Conclusion
We investigate a series of critical yet unexplored questions regarding the effectiveness and feasibility of token-precision trade-off in KV cache compression. Through comprehensive experiments, we demonstrate that storing more tokens in the KV cache with lower precisioncan significantly enhance the long-context performance of LLMs, and demonstrating robust performance across diverse input lengths, downstream tasks, with particularly significant gains in retrieval tasks. Moreover, we find quantized pruning demonstrates strong feasibility across different KV pruning methods, quantization strategies, and model scales. Our analysis sheds light on the token-precision trade-off of KV cache memory optimization, offering valuable insights into designing more efficient compression strategies. We hope this work deepens our understanding of KV cache compression and inspires future research.
@@ Appendix A Datasets
LongBench (Bai et al., 2024) includes 17 datasets covering 6 categories of tasks, which can be divided into single-document QA (Dasigi et al., 2021; Kočiskỳ et al., 2018), multi-document QA (Yang et al., 2018; Ho et al., 2020), summarization (Huang et al., 2021; Fabbri et al., 2019; Zhong et al., 2021), few-shot learning (Gliwa et al., 2019; Joshi et al., 2017; Li and Roth, 2002), synthetic, and code generation (Guo et al., 2023; Liu et al., 2023). LongBench features an average input length ranging from 1,235 to 18,409 tokens. For inputs exceeding the model’s context window length, we split the data and only take the beginning and end segments of the input to fill the context window length. For Input Lengths Experiment, we select datasets with sufficient data to cover three input length ranges: (<4k, 4k 8k, and >8k). These datasets include MultiFieldQA-en, 2WikiMultihopQA, GovReport, TREC, TriviaQA, SAMSum, and RepoBench-P, representing a variety of task types. We refer to the three subsets as LB-4k, LB-8k, and LB-16k, respectively.
@@
Needle-in-a-Haystack(NIAH) (Kamradt, 2023) is a challenging pressure test designed to assess the ability of models to accurate identify and retrieve relevant information from lengthy context. NIAH randomly inserts key information into an arbitrary position within a long essay. In our setup, we use PaulGrahamEssays as the haystack and the sentence "The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day." as the needle, which is the default setting of NIAH. We vary the essay length from 1,000 tokens up to the models’ context window limits, increasing by 100 tokens per step for Llama-series models and 400 tokens per step for Mistral.
@@
Figure 7: The results of Layer-Wise Quantized Pruning on Mistral-7B-v0.2-Instruct.
@@ Appendix B Experiment Setup
We retain the last 32 tokens for StreamingLLM (Xiao et al., 2023), H2O (Zhang et al., 2024b), and SnapKV (Li et al., 2024), while keeping 8 tokens for PyramidKV (Cai et al., 2024), as recommended in the corresponding paper (Cai et al., 2024). For other settings, we adopt the default configurations from the PyramidKV codebase.
@@ Appendix C More results in Experiments
We alse report the results for Mistral-v0.2 in Figure 7, the results are similiar to Llama-3.

### v2

@@ 6.2 Scaling Effect on Quantized Pruning
Figure 3: Scaling effect on Llama family models, with PyramidKV-based quantized pruning. All models are under 1/64 KV cache budget.
To investigate the impact of model scaling on quantized pruning, we conducted experiments on four models from the Llama series: Llama3-70B, Llama3-8B, Llama3.2-3B and Llama3.2-1B. Since the Llama3 series does not include 1B and 3B versions, we used the Llama3.2 series as substitutes. However, it is important to note that the performance of Llama3.2-3B-Base is quite similar to Llama3-8B-Base. For both the Base models and Instruct models, we evaluated their performance on LongBench under 1/64 KV cache budgets. To further validate the conclusion, we also experimented with a 1/16 KV cache budget, and the results are presented in the Appendix C.
As shown in Figure 3, we observe that quantized pruning consistently achieves better performance across all scaling levels. Notably, the performance gap between quantized pruning and standalone KV pruning methods remains relatively stable across different model scales. Interestingly, as the model scale increases, the performance gap between 2× tokens with 8-bit precision and 4× tokens with 4-bit precision becomes smaller. This may be because deeper models encode more redundant information within each token, allowing a relatively smaller number of tokens to retain sufficient information. For Base models, although the performance improvement from scaling is smaller compared to Instruct models, quantized pruning still provides a noticeable performance boost.
@@ 6.3 Ablation on Quantization Strategies
While there has been extensive research on strategies for KV quantization, it remains unclear whether existing quantization strategies remain effective when integrated with KV pruning methods. In this section, we aim to investigate the impact of KV quantization strategies on quantized pruning. We also conducted an analysis of the effect of quantization group size on the quantized pruning, with further results available in Appendix C.
Quantization methods
We explore the methods in FlexGen (Sheng et al., 2023), KIVI (Liu et al., 2024c), and KVQuant (Hooper et al., 2024). To elaborate, for the FlexGen methods, KV quantization is applied to both the key and value caches along the token dimension, grouping every 64 elements without filtering outlier numbers. We modify the FlexGen by (1) filtering 1% of outlier numbers in both the key and value cache, as mentioned in KVQuant (2) quantizing the key along the channel dimension, as in KIVI and (3) combining (1) and (2). These correspond to the results labeled as FlexGen+Outlier 1%, KIVI, and KIVI+Outlier 1% in the Figure 4.
We can observe that none of the quantization strategies show significant performance degradation when combined with KV pruning methods, demonstrating the relative stability of quantized pruning. The KIVI method consistently outperforms FlexGen across various models and KV pruning methods. The improvement is particularly pronounced for PyramidKV on the Mistral model, underscoring the significance of quantizing key states along the channel dimension. Filtering 1% of outlier numbers proves effective for the FlexGen strategy but yields limited improvements for KIVI. It shows some benefit on Llama models but results in negligible gains on the Mistral models.
Overall, KIVI demonstrates strong performance when integrated with KV pruning methods, while other KV quantization strategies also maintain good results, highlighting the stability of quantized pruning.
6.4 Exploration on Layer-Wise Quantized Pruning
Inspired by the observation that different layers may have varying requirements for the number of tokens in PyramidKV (Cai et al., 2024) and PyramidInfer (Yang et al., 2024a), we further investigate whether the demands for precision and preserved tokens are consistent across layers. To explore this, we use the best-performing configuration from previous experiments, 4-bit precision with 4×4$\times$4 × tokens, as the baseline and compare it against layer-wise configurations adopting 8-bit precision with 2×2$\times$2 × tokens and 16-bit precision with 1×1$\times$1 × tokens. Using SnapKV as the KV pruning method, we present the results for Llama-3 under 1/64 KV cache budget in the Figure 5. The x-axis shows the modified layers range, while the y-axis shows the relative change to the baseline (4-bit precision with 4×4$\times$4 × tokens) on LongBench and RULER-4k. Furthermore, we present the results for Llama-3 evaluated with a 1/16 KV cache budget and Mistral-v0.2 evaluated with 1/64 and 1/16 KV cache budgets. These detailed results are available in the Appendix C.
@@ 6.4 Exploration on Layer-Wise Quantized Pruning
Figure 5: The results of layer-wise quantized pruning on Llama-3-8B-Instruct, with SnapKV as pruning method. We use 4×4$\times$4 × KV token 4-bit as baseline and report the relative change. Configurations are modified every 4 layers for the initial and final layers, while intermediate layers are reconfigured every 8 layers.
@@
It is evident that for most layers, transitioning from 4×4$\times$4 × tokens with 4-bit precision to higher precision and fewer tokens results in a performance decline under constrained KV cache budgets. Specifically, the shift to 8-bit shows a relatively minor performance drop, whereas moving to 16-bit with fewer preserved tokens leads to a more significant decrease. These layers-wise trade-off conclusions are consistent with our experiments before.
@@
Notably, modifying intermediate layers causes a drastic performance decline, while adjustments made at the initial and final layers result in comparatively smaller performance reductions. This effect is especially pronounced in retrieval-related tasks such as RULER-4k, where significant performance differences are observed. On LongBench, changes are less significant, with a notable performance drop only observed at 16-bit precision. These findings highlight that, under the same memory budget, preserving more tokens with lower precision in intermediate layers is crucial for the performance, while the token-precision trade-off in the initial and final layers exerts a more balanced influence.
@@ 7 Conclusion
We investigate a series of critical yet unexplored questions regarding the effectiveness and feasibility of token-precision trade-off in KV cache compression. Through comprehensive experiments, we demonstrate that storing more tokens in the KV cache with lower precision can significantly enhance the long-context performance of LLMs, and demonstrating robust performance across diverse input lengths, downstream tasks, with particularly significant gains in retrieval tasks. Moreover, we show that quantized pruning demonstrates strong feasibility across different KV pruning methods, quantization strategies, and model scales. Our analysis sheds light on the token-precision trade-off of KV cache memory optimization, offering valuable insights into designing more efficient compression strategies. We hope this work deepens the understanding of KV cache compression and inspires future research.
@@ Appendix A Datasets
LongBench (Bai et al., 2024) includes 17 datasets covering 6 categories of tasks, which can be divided into single-document QA (Dasigi et al., 2021; Kočiskỳ et al., 2018), multi-document QA (Yang et al., 2018; Ho et al., 2020), summarization (Huang et al., 2021; Fabbri et al., 2019; Zhong et al., 2021), few-shot learning (Gliwa et al., 2019; Joshi et al., 2017; Li and Roth, 2002), synthetic, and code generation (Guo et al., 2023; Liu et al., 2023). LongBench features an average input length ranging from 1,235 to 18,409 tokens. For inputs exceeding the model’s context window length(8k for Llama-3-8B-Instruct (Dubey et al., 2024), we split the data and only take the beginning and end segments of the input to fill the context window length. Additionally, we reserve sufficient space for newly generated tokens based on the specific type of sub-dataset. For evaluate the impact of input lengths, we select datasets with sufficient data to cover three input length ranges: (<4k, 4k 8k, and >8k). These datasets include MultiFieldQA-en, 2WikiMultihopQA, GovReport, TREC, TriviaQA, SAMSum, and RepoBench-P, representing a variety of task types. We refer to the three subsets as LB-4k, LB-8k, and LB-16k, respectively.
@@
Needle-in-a-Haystack(NIAH) (Kamradt, 2023) is a challenging pressure test designed to assess the ability of models to accurate identify and retrieve relevant information from lengthy context. NIAH randomly inserts key information into an arbitrary position within a long essay. In our setup, we use PaulGrahamEssays as the haystack and the sentence "The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day." as the needle, which is the default setting of NIAH. We vary the essay length from 1,000 tokens up to the models’ context window limits, increasing by 100 tokens per step for Llama-series models and 400 tokens per step for Mistral. The results are reported as the average score across all tests.
Figure 6: Scaling effect on Llama family models, with PyramidKV-based quantized pruning. All models are under 1/16 KV cache budget.
@@ Appendix B Experiment Setup
We retain the last 32 tokens for StreamingLLM (Xiao et al., 2023), H2O (Zhang et al., 2024b), and SnapKV (Li et al., 2024), while keeping 8 tokens for PyramidKV (Cai et al., 2024), Ada-KV (Feng et al., 2024) and HeadKV (Fu et al., 2024), as recommended in the corresponding paper (Cai et al., 2024; Fu et al., 2024). For other settings, we adopt the default configurations from their papers.
@@ Appendix C More results in Experiments
Group Size
We analyze the impact of group size during KV quantization. We employ SnapKV and PyramidKV to retain 512 tokens with 4-bit KIVI quantization and report the performance variations when the group sizes were set to 32, 64 and 128. As shown in Table 3, smaller group sizes lead to performance improvements at the cost of higher memory usage. Reducing the group size from 128 to 64 resulted in a notable improvement, but further decreasing it from 64 to 32 yielded minimal gains for the Mistral model. Therefore, we set the default quantization group size to 64 to balance performance and memory usage in our experiments.
Model Method Group Size
32 64 128
Llama-3 SnapKV 40.4 39.6 38.9
PyramidKV 40.3 39.6 38.9
Mistral-v0.2 SnapKV 40.4 40.3 40.1
PyramidKV 40.3 40.5 40.0
Table 3: The impact of group size for quantized pruning on LongBench.
Scaling Effect on Quantized Pruning
We validate our findings by measuring performance at a budget of 1/16, with the results shown in Figure 6. This corroborates the conclusion reported in the main text using a budget of 1/64. When the KV cache budget is relative small to 1/64 , the performance improvement brought by quantized pruning is higher compared to 1/16 KV cache budget, which aligns with the conclusions we observed earlier in Q2.
Figure 7: The results of Layer-Wise Quantized Pruning on Llama-3-8B-Instruct, with SnapKV as pruning method. KV cache budget=1/16.
Figure 8: The results of Layer-Wise Quantized Pruning on Mistral-7B-v0.2-Instruct, with SnapKV as pruning method. KV cache budget=1/64.
Figure 9: The results of Layer-Wise Quantized Pruning on Mistral-7B-v0.2-Instruct, with SnapKV as pruning method. KV cache budget=1/16.
@@
We further validated our findings by measuring performance of Llama-3 at a budget of 1/16, with the results shown in Figure 7. We alse report the results for Mistral-v0.2 in Figure 8 and Figure 9, we can see the layer-wise results are similiar to Llama-3. Modifying intermediate layers causes a drastic performance decline, while adjustments made at the initial and final layers result in comparatively smaller performance reductions. This effect is especially pronounced in retrieval-related tasks such as RULER-4k, where significant performance differences are observed. On LongBench, changes are less significant, with a notable performance drop only observed at 16-bit precision.